### v1

# The Atacama Cosmology Telescope: The Development of Machine Learning Tools for Detecting Millimeter Sources in Timestream Pre-processing

Abstract
We present a new machine learning algorithm for classifying short-duration features in raw time ordered data (TODs) of cosmic microwave background survey observations. The algorithm, specifically designed for the Atacama Cosmology Telescope (ACT), works in conjunction with the previous TOD preprocessing techniques that employ statistical thresholding to indiscriminately remove all large spikes in the data, whether they are due to noise features, cosmic rays, or true astrophysical sources in a process called “data cuts”. This has the undesirable effect of excising real astrophysical sources, including transients, from the data. The machine learning algorithm demonstrated in this work uses the output from these data cuts and is able to differentiate between electronic noise, cosmic rays, and point sources, enabling the removal of undesired signals while retaining true astrophysical signals during TOD pre-processing. We achieve an overall accuracy of 90% in categorizing data spikes of different origin and, importantly, 94% for identifying those caused by astrophysical sources. Our algorithm also measures the amplitude of any detected source seen more than once and produces a sub-minute to minute light curve, providing information on its short timescale variability. This automated algorithm for source detection and amplitude estimation will be particularly useful for upcoming surveys with large data volumes, such as the Simons Observatory.
1 Introduction
Ground-based microwave astronomy has experienced a revolution in the past two decades, thanks to the deployment of large aperture telescopes that have thousands of detectors, increased their resolution, and advances in detector and read-out technology that have drastically improved their sensitivity while also allowing for simultaneous observation of multiple wavelength bands. For example, the Atacama Cosmology Telescope (ACT; Swetz et al. 43, Thornton et al. 45, Henderson et al. 25) observed the cosmic microwave background (CMB) with 3000 detectors across up to three frequency bands, covering six bands from 30 to 270 GHz during its operation from 2007 to 2022. ACT will be succeeded by the Simons Observatory (SO), which is initially set to deploy over 60,000 detectors split between large- and small-aperture telescopes across six frequency bands from 30 to 280 GHz 3, 2. During observations, detectors, such as transition edge sensors (TES), record the incident power at a cadence of a few hundred hertz as the telescope scans in azimuth at a constant elevation. In the ACT survey, the data timestreams of all detectors are packaged into files known as “time-ordered data streams” (TODs), each containing roughly 11 minutes worth of data. Using knowledge of the telescope pointing, which is recorded during data collection, the TODs are later processed into microwave sky maps using maximum likelihood map-making 19, 17, 4.
In addition to using the CMB directly as a probe of cosmological physics, an emerging use of CMB experiments is the study of the time-varying microwave sky. Wide-area surveys contain thousands of radio sources; the majority of them are blazars that emit microwave light that changes in flux over days/weeks. The South Pole Telescope (SPT) collaboration 15 published a pilot study of one blazar (PKS 2326-502) observed in the microwave bands 28. Surveys like ACT and SO will provide much longer baselines of observations of such objects: the field of mm-transient studies is expected to grow significantly in the coming decade. Similarly in the realm of planetary science, ACT and SPT have also characterized the millimeter flux of asteroids 16, 39, and ACT constrained the existence of a “Planet 9” in our Solar System 36. Finally, ACT and SPT have successfully measured dozens of transients in their data that are mainly flaring stars, with the majority consisting of short-duration (minutes- to hours-long) stellar flares 49, 37, 31, 22, 44, 12. Additionally, as part of a blind transient search, ACT has detected classical nova YZ Ret during its outburst, making it only the second millimeter observation of this kind 12. ACT also published a targeted search for extragalactic transients that included gamma ray bursts (GRBs), tidal disruption events (TDEs), and supernovae (SN), placing upper-limits on the microwave flux of the target objects 26. More sensitive data from the upcoming SPT, SO, and CMB-S4 1 surveys will produce detections in the millimeter of extragalactic transients, specifically GRBs, which will provide valuable information on the emission processes of these powerful explosions 20, 2.
CMB surveys are designed to measure the power of faint microwave radiation imprinted on all angular scales across the sky. However, the data collected by ground-based CMB telescope like ACT are dominated by microwave emission from water vapor in the atmosphere and are additionally contaminated from, e.g., instrumental noise and cosmic rays colliding with the detectors. While instrumental and atmospheric noise can be modeled in the microwave map-making process to produce unbiased, maximum likelihood maps 19, the mapmaker is specifically vulnerable to biases from short, high-amplitude, non-Gaussian data bursts in the TOD, and therefore depends on these bursts being excised from the TOD before map-making. Such transient events can be caused by sporadic pathologies from the electronic readout system, cosmic rays, electromagnetic interference, radio frequency interference modulated by telescope motion, ionizing radiation, digitization artifacts, etc 9, 8, 33. Cosmic rays, for example, sometimes produce rapid spikes in the TODs of single detectors and may sometimes also warm up a region of the detector polarization array (PA) near the impact, thereby inducing a signal spike in multiple detectors (see 33 for a study on cosmic rays for Planck HFI). In this paper, we refer to all of these transient phenomena as “glitches”. Bright astrophysical point sources can also cause high-amplitude, non-Gaussian spikes in the TOD and may be mistaken for a glitch. This may be avoided if the coordinate of the source is known before pre-processing. However, this also means that any point source or astrophysical transient that is not already known in advance of map-making can be removed from the data set entirely. We note that while the specific phenomena we describe here are commonly encountered in ACT timestreams, other transient sources may manifest differently in different telescopes depending on their readout schemes and observing conditions.
It is therefore imperative when pre-processing TODs from CMB telescopes to excise as many glitches as possible while still retaining as many true astrophysical transients as feasible. As the volume of data observed by current and future experiments increases, traditional techniques that rely on human inspection of a significant fraction of candidate glitches in order to tune statistically-based thresholding algorithms will rapidly become unwieldy. Improvements in the data-excision algorithms are needed, both to better automate the process and to avoid cutting real signals. In this paper, we describe a new machine learning method that achieves these objectives utilizing the outputs of the current glitch finder.
After reviewing in Section 2 the “cuts process” used by ACT to cut out potential glitches from the data, we describe our new approach in Sections 3 and 4, and introduce our supervised machine learning algorithm in Section 4. We explain the summary statistics derived from the TODs that are fed into the algorithm in Section 4 and how we prepare our training and test sets in Section 5. We test the classification method on simulated stellar flares which are discussed in Section 6, summarize and discuss our results in Section 7, and present our conclusions and suggestions for future work in Section 8.
2 Current State of ACT Data Cuts
Before raw ACT TODs are processed into CMB maps by the map-making pipeline, portions of the TODs suspected to be glitches or contaminated with noise are excised from the data in a process known as “data cuts”. The data cuts algorithm in ACT is described in 4 and in detail in 19; however, we summarize some key steps and properties of the cuts algorithm below.
Firstly, the data cuts algorithm can discard entire TODs based on the level of precipitable water vapor in the atmosphere during the TOD, or the number of faulty detectors operating over that time period. Additionally, we flag individual detectors within each TOD based on various statistical metrics computed on the TOD. A detector may become faulty for various reasons, such as saturation from optical and electric power, an instability caused by insufficient heat capacity, slow time constants (which describes how quickly the detector changes in response to a signal) from fabrication and wiring problems, or failures in the detector readout system. Current methods for producing individual detector cuts utilize a set of parameters to identify and exclude such detectors with anomalous behavior. These anomalies often manifest as abnormally high or low noise levels, poor optical coupling, or excessive high-frequency, non-Gaussian noise above around 10 Hz, where photon shot noise with a white spectrum is expected to dominate. Detectors that have outlier values of the statistics for these parameters are flagged and rejected from map-making.
Moreover, short, high signal-to-noise (SNR) spikes on millisecond (ms) time scales are also prevalent in the detector timestreams, caused by glitches in the readout system and cosmic rays, among other factors. ACT employs a glitch finder that combines high-pass and Gaussian filters to identify samples affected by glitches, defined as those with in the high-pass-filtered TOD. This method is effective because ACT TODs are predominantly noise-dominated. When a glitch is flagged, a buffer of 200 time samples on either end of the glitch is identified, and all samples from the detector within this buffer interval are rejected from the map-making process.
While this glitch finder is effective at removing glitches due to non-astrophysical signals, it also excludes bright astronomical sources that appear as spikes in the timestream as the telescope scans across them. Given the ACT scanning speed () and beam size arcminute at 150 GHz, astrophysical source spikes typically last around 20–30 ms. In the standard map-making pipeline, a catalog of bright astrophysical sources is curated and passed to the glitch finder; samples in the TOD that lie within of a known source from the catalog are masked before running the map-maker. This approach ensures that known bright astrophysical sources will not be cut, but it necessitates multi-pass TOD processing and, more importantly, may fail when a previously unknown source of transient nature appears in the observing field, such as those reported in Naess et al. 37 and Li et al. 31. If these signals are bright enough, they might be erroneously removed by the glitch finder, resulting in missed opportunities to discover new sources of astrophysical interest.
Different types of glitches typically manifest differently in the timestream and across the focal plane. For instance, cosmic rays often simultaneously affect detectors in close proximity to each other. In contrast, a signal from a bright astrophysical point source is extended on the focal plane given the finite beam-shape. Point source signals are thus extended in the timestream as well and appear at different times for different detectors depending on their locations on the focal plane. As the telescope scans the sky at a constant elevation, detectors that see the same source fall along a constant-elevation line on the focal plane during each scan. The distinct spatial and temporal characteristics of different glitches can in principle be used to differentiate them. The spatio-temporal information of the TOD data has also been successfully used to model the bulk motion of the atmosphere 34.
In this work, we demonstrate that by combining both spatial and temporal information for different classes of glitches, it is possible to effectively distinguish between them with the assistance of a novel machine learning model. This approach provides insights into the physical origins of each glitch, and offers a promising method for preserving and identifying astrophysically interesting transient signals in the TODs in a fully automated method without the need for visual inspection by human experts. This property of the glitch classification will have growing importance as the data volume scales up in the coming decade with surveys like SO.
3 Data Preparation
Within the methodology we present here, the first step is to calibrate and perform data cuts on each TOD, following the same steps to identify and excise glitches as outlined in Section 2. Since each physical event that triggers glitches in a TOD may affect multiple detectors at slightly different times, we group together glitches that occur in different detectors within 200 time samples of each other (ACT has a sampling rate ranging from 300 - 400 Hz depending on the detector array). For example, if glitch 1 and glitch 2 occur within 200 samples of each other, and glitch 3 and glitch 2 occur within 200 samples of each other but glitch 1 does not, all three glitches are grouped together. We extract the samples from the impacted detectors within this joint interval and call that subset of the timestream a snippet. Snippets thus group together samples from multiple detectors that are potentially affected by the same physical event, and can be analyzed to reveal the spatial and temporal structure of that physical event.
Figure 1 shows snippets for the four categories we define in this work to generally summarize the phenomena we see in the glitches for classification: point sources (PS), point sources with another coincident glitch (abbreviated as PS+), cosmic rays (CR), and electronic glitches (EG). In the figure, the example snippet for the PS+ contains a point source together with a cosmic ray, but this category can also include an electronic glitch rather than a cosmic ray. There is a large variety of electronic glitches that are observed in our data (all labeled as EG): for instance, in the bottom panel of Figure 1, a diagonal “striping” is seen on the focal plane. These detectors share a common read-out line in the time-domain multiplexed system of our Multi-Channel Electronics (MCE) readout system 10. This particular electronic glitch has affected the whole read-out line. For our machine learning classification of these categories, we characterize both the TODs and layout of the affected detectors across the focal plane using the summary statistics described in the following section.
4 Description of Classification Features and Hyperparameters
We use a random forest machine learning algorithm, which can be thought of as a collection of “decision trees” that classify data based on the input features 13. In order to differentiate the glitches produced by the current ACT glitch finder, we extract seven summary statistics per snippet from the timeseries and detector positions as features for the random forest, which can be further grouped into five distinct subtypes:
1.
the number of detectors affected by the glitch (Number of Detectors; Section 4.1);
2.
summaries of the distribution of these detectors across the focal plane (Y and X Extent Ratio, Normalized Y Maximum, and Normalized Y Maximum Within 0.1; Section 4.2);
3.
the correlation of the signal (Correlation; Section 4.3);
4.
the time lag between detectors (Time-lag; Section 4.4);
5.
and the number of peaks in the TODs (Number of Peaks; Section 4.5).
Figure 2 shows the distributions of the seven statistics, described in the subsections that follow, falling in each of the five types enumerated above.
4.1 Number of Detectors
We consider the number of detectors affected by a glitch as one of our summary statistics. Generally for a point source we expect (but generally fewer than 100) detectors to be affected depending on where the point source crosses the focal plane as the telescope scans in azimuth, as only a strip of detectors will observe the source. A cosmic ray will generally only affect a small cluster of detectors; but they can affect large portions of the focal plane, depending on the energy of the cosmic ray and if multiple cosmic rays hit the focal plane at the same time. The largest spread in the number of affected detectors occurs for the EG category, as it includes a large variety of glitches. Many electronic glitches will only affect detectors, however there are instances, for example large-area heat-ups, where almost every detector on the focal plane is affected. Finally, when a point source occurs coincident with another glitch, we expect a slightly larger number of affected detectors than for the PS category, since there is another glitch occurring as well.
4.2 Detector Distributions Across the Focal Plane
Multiple summary statistics are used to help characterize the affected detector distributions across the focal plane. Each detector has a unique - and -coordinate in degrees on the focal plane. In order to compute useful statistics, the focal plane distributions are collapsed along the - and -axes to create histograms summarizing these distributions. We would expect that stationary astrophysical sources to affect detectors along the scan direction which corresponds to the -axis and the -coordinate distribution to fall within a few beams. Example focal planes for the glitch categories can be seen in Figure 1 and example histograms for a PS and a PS+ are shown in Figure 3.
4.2.1 The and Extent Ratio
One of the statistics that can be derived from the focal plane is the ratio of the ranges in the and directions, which is computed as:
where are the maximum and minimum values of the distributions of the affected detectors respectively.
Since many detectors will observe a PS across the focal plane in a horizontal line as the telescope scans in azimuth, we expect a large spread in the coordinate but only a small range of affected detectors in the direction, since the celestial sphere barely rotates during a single scan. Therefore, there should generally be a small ratio of extent compared to the extent/range for a point source. A cosmic ray will hit a cluster of detectors, which results in a similar spread in the and directions, and we therefore generally expect the ratio to have a value close to unity for CR objects. However, this is not always the case. For example, if multiple cosmic rays hit the focal plane at once, the ratio will deviate from unity, but the “ and Extent Ratio” will typically remain larger than that of a point source. The expected extent ratio also varies quite significantly for the EG category depending on the type of electronic glitch, however the ratio is generally larger than that for a point source. For a PS+, we expect a variety of deviations that differ from the extent ratio expected for a point source alone, depending on the location of the other glitch with respect to the location of the source on the focal plane. If the glitch occurs near the source, we similarly expect a small ratio. Conversely, if the glitch occurs far away on the focal plane, the ratio will be closer to one as it will have a large range and the range is generally close to the full extent of the focal plane for a point source.
4.2.2 Normalized Maximum
To derive this statistic, we compute the histogram of the distribution of the detectors in the direction of the focal plane for a given glitch. The histogram bin with the largest number of detectors and the adjacent bins are added together and then normalized by the total number of detectors present in the glitch. We compute the histogram of detectors with 10 bins, resulting in the width of an individual bin in the histogram being small when the detectors are clustered and larger if there is spread in positions. Figure 3 shows examples for both the PS and PS+ categories, where the detectors being considered are shown as coral compared to all the detectors in the snippet shown in purple. Note that in both examples, there is only one bin beside the maximum bin, thus only two bins are used for the “Normalized Maximum” statistic. For the PS+, this is because the maximum bin is the first bin, whereas for the PS it is due to the detectors that would have been in the bin to the right of the maximum being cut during the detectors cuts which happen before the glitch cuts. The PS object, has only 27% of the detectors within the maximum and adjacent bins, compared to 59% of detectors within the maximum and adjacent bins for the PS+ object.
This statistic is most useful for recognizing a PS+ object, as there tend to be numerous detectors that observe the point source itself compared to those that see only the coincident glitch. Conversely, because point sources and cosmic rays are typically tightly clustered and have less scatter in the direction, a 10-bin histogram of the detectors has a smaller individual bin width and generally does not contain a large fraction of the affected detectors. Objects classified as EG have a variety of distributions across the focal plane and are thus not well characterized by this statistic.
4.2.3 Normalized Maximum Within 0.1
In general, we expect the spread of affected detectors in the direction to fall within 0.1, or a few beams, for an astrophysical source. This is not true for a general electronic glitch or cosmic ray, which makes this metric useful for identifying astrophysical sources. We compute the fraction of detectors in the direction of the focal plane that are within 0.1 of the histogram bin that contains the maximum number of detectors. Figure 3 shows examples for both the PS and PS+ categories for both the “Normalized Maximum” and “Normalized Maximum Within 0.1” to compare these statics.
Similarly to the previously defined statistic, we expect a small range of detectors to observe a point source as the telescope scans over the source. Therefore, the number of detectors within 0.1 of the position of the maximum histogram bin should be close to the total number of detectors in the snippet (and therefore our statistic normalized over the total number of detectors should be very close to unity, as shown in Figures 2 and 3). This makes the statistic very useful for determining if a glitch is a point source. A cosmic ray will hit a small cluster of detectors, resulting in a small spread in the direction. We still expect the fraction of the total number of detectors within 0.1 of the position to be close to unity, but there can be deviations from this if multiple cosmic rays hit the focal plane at once, and they are still generally smaller than for a point source. The expected distribution of detectors varies quite significantly depending on the type of electronic glitch is occurring, but we find that the statistic is generally smaller for electronic glitches than for the PS class. For a PS+ object, the value of this statistic will differ from what is seen for an isolated source depending once again on where the additional glitch occurs on the focal plane relative to the source location: if the glitch occurs near the source, we similarly expect many detectors to be included, leading to a large fraction of the detectors to be within 0.1 of the maximum. Conversely, if the glitch occurs far away on the focal plane, this will result in a smaller fraction of detectors within 0.1 of the maximum and a smaller value for our statistic. This statistical metric is therefore useful in identifying PS+ objects where the coincident glitches are located elsewhere in the detector array and thus would be able to be easily separated in the snippets from the point source and for further analysis of the source.
4.3 Correlation
In order to determine how the incident flux is correlated across detectors (which can help distinguish between the PS and CR classes), we compute the Pearson correlation coefficient of each detector pair ordered by their position across the focal plane. We then use the mean of the absolute value of these correlation coefficients from all detector pairs in the snippet as a feature in our classification algorithm. Given that detectors observe the point source sequentially as the telescope scans across the sky, we only expect a large value of this correlation coefficient between detectors that are either close to or beside the main detectors on the focal plane that observe the source. This scenario leads to a block diagonal correlation matrix of high values across the array when the detectors are ordered by their position (see Figure 4 for an illustration) but a low overall degree of correlation across all detectors. For cosmic rays, we expect a high level of correlation across the array, because the detectors are all affected by the cosmic ray at the same time, shown in Figure 4. For electronic glitches that cause a strong signal that affects multiple detectors simultaneously, we expect a large value of the correlation coefficient during, e.g., a large-area heat-up. Conversely, if the glitch is similar to a random noise fluctuation, we would expect a low value of the correlation coefficient. For PS+ objects, if the additional glitch is a cosmic ray, we expect a slightly higher correlation than would be observed for a point source, as the cosmic ray portion of the TODs for each detector will overlap while affected by the cosmic ray. If the additional glitch is of electronic or other origin, the effect varies, but it generally still increases the level of correlation, as shown in Figure 2.
4.4 Time-lag
While the absolute value of the correlation described above indicates the spatial correlation between detectors that see a given glitch, the temporal correlation (or time-lag) is also a key statistic that we use in classification. The time-lag between detectors is the shift in time that maximizes the cross-correlation between the two detector TODs over the snippet. For two detectors and , it is defined following 35,
where DFT is the discrete Fourier transform, the overbar denotes the complex conjugate, and is the timestream of the detector as a function of time. We take the absolute value, since the sign of the shift (backwards or forwards) does not contain information relevant to the type of the signal. We compute this time-lag for all detector pairs in the snippet, and then use the mean of the absolute values to summarize the glitch.
For a point source, each detector observes the source sequentially as the telescope scans across the sky. We therefore expect a time-lag that is equal to the separation in azimuth of the two detectors divided by the scanning speed. In practice, we can get different values of the time-lag based on how noisy the TOD is, especially for low amplitude sources. This quantitative number can be used in future work to differentiate an astrophysical point source from a moving source, such as a satellite. Conversely, we expect a time-lag close to zero for cosmic rays, since the cosmic ray timescale is fast compared to the readout rate, which means that the digitized signatures reside essentially on top of each other in the detector timestream. This can be seen in Figure 5, where the time-lag for detector pairs, ordered by their position, are shown for a cosmic ray on the top and a point source on the bottom. As expected, the time-lag is zero for almost every detector pair for the cosmic ray. For the EG category, it will cause varied time-lags depending on the nature of the glitch. Generally, something such as a focal plane heating event or another electronic glitch that affects all the detectors at the same time will result in a zero time-lag as the detectors experience the effect at the same time. Finally, for the PS+ class, if the other glitch is a cosmic ray, we expect a slightly smaller time-lag than just a point source as part of the TODs will be overlapping with a zero time-lag. If the other glitch is of an electronic or other origin the effect varies, but generally speaking it increases the time-lag, as we show in Figure 2.
4.5 Number of Peaks
In a given snippet, some glitches, in particular glitches from PS objects and CR objects, will appear as “peaks” in the TOD, and so the number of signal peaks at unique times is used as one of our summary statistics, as it is particularly useful in discerning a cosmic ray from an electronic glitch. In order to compute the number of peaks in a snippet, we first smooth the TODs for the snippet using a boxcar smoothing kernel of width three time samples to remove small fluctuations due to random noise. We then combine the TODs from each detector into a single timestream , as:
where and indicate the maximum and mean of the all the samples at a given time , respectively, and is the standard deviation of all detectors and times in the snippets. Using this combined TOD, we use the Python scipy.signal.find_peaks function to find the number and locations of the peaks for the TOD snippet. Examples for both a cosmic ray and a point source are shown in Figure 6. PS objects and combined PS+ objects generally have peaks, whereas most electronic glitches display fewer than five peaks, as can be seen in Figure 2. Signals from cosmic rays generally only have one peak unless multiple cosmic rays hit simultaneously, which makes this statistic useful for breaking the degeneracy between them and electronic glitches.
4.6 Importance of Statistics
When using the various statistics described above, it is important to determine which of the features are more instructive in the classification step, to ensure that we understand how the classifier weights them. Low-importance features can potentially be removed from a feature set, and high-importance features can be examined to identify the physics behind why the features lead to more accurate classifications. It also aids in our understanding of misclassified glitches; see Section 7.1 for more details.
We compute the relative importance of the various features described above to train our random forest classifier using the “permutation feature importance” method 13. This method randomly shuffles the values of each feature and computes the change in the model performance for those values. The impact of this on the performance of the classifier indicates the importance of each feature in determining the classifications. We randomly shuffle the features 50 times each to reduce the variability of single permutations, and show the distributions of the feature importance for each statistic in Figure 7. A larger decrease in the accuracy score indicates a higher importance of the feature, as it shows that changing the values of those statistics would have a more significant impact on the classification results (and vice versa). For our model, the focal plane statistics tend to have the largest impact on the accuracy score. This is discussed further in section 7.1.
4.7 Optimizing the Machine Learning Algorithm
We explored the performance of the random forest as a function of the number of trees and the “max depth” (which is the maximum number of times each decision tree is allowed to split) of the forest classifier. The number of decision trees used was varied between 25 and 175 (preliminary tests outside that range showed minimal impact in performance) and the max depth varied within an initial range of 2 to 32. The performance was assessed using the overall accuracy of the classifier, the overall out-of-bag (OOB) score, and F-1 score which is the harmonic mean of the precision and recall (see Section 7.1 for more information on our metrics) for the four glitch categories. The random forest approach uses bootstrap aggregation to train each tree, a process whereby the training set for each decision tree is produced by sampling from the original training set with replacement. This avoids overfitting of the algorithm to the training data, which can lead to a model that is does not perform well on new (test) data. The ability of a method like random forests to make predictions on data that were not included in the training set for each individual decision tree within the ensemble of trees is characterized by the OOB score 24. The performance metrics discussed above plateau and do not improve when the forest reaches a size of 50 trees and a max depth value of 15; therefore, we fixed those hyperparameters of the random forest method to these values.
5 Training and Test Set Preparation
5.1 Labeling of Objects in the Training and Test Sets
Conventionally, the creation of training and test data proceeds via human (manual) labeling. In order to create a large enough training set for the random forest classifier, we needed an efficient method to facilitate manual labeling of the snippets, via a graphical user interface (GUI).
We use the Python-driven platform developed by the Zooniverse collaboration111www.zooniverse.org to create the labeling GUI. This Zooniverse platform includes a well-established data pipeline and a user-friendly interface, as can be seen in Figure 8. The platform we employ was designed specifically for use by members of the ACT collaboration who can simultaneously label glitches stored on a common computing facility, to assist in both training and verification of the algorithm.
To facilitate classifications by non-expert users, we created a user tutorial specific to ACT glitch data, with examples related to the key features of each type of glitch and notes on pitfalls in classifying glitches.222We separately included instructions on setting up a Zooniverse project from scratch, see online documentation here.
As shown in Figure 8, the user is shown a figure of the detector array, where the red detectors are those affected by the glitch. The user is also shown a figure of the TODs from all affected detectors superposed as a snippet, and is provided with five options for possible glitch labels (the four labels described in Section 3 and an additional “None of the above” label).
The test and training data are then drawn from one larger sample of objects labeled in the manner described above. We describe the procedure for the separating this data set into a training set and a test set below.
5.2 Forming the Training and Test Sets
Once we have obtained a representative sample for training and testing our classifier, we need to ensure we are robust to changes in the snippets from different years (observation seasons), observational frequency bands, detector array configurations, etc., which requires further refinement of our test/training split. In order to ensure we are robust to these changes, we use an active learning prescription, in particular the modAL 18 learning framework. Active learning is a machine learning approach where unlabeled objects are selected for manual classification/labeling depending on specific criteria of the object (for example the classification ambiguity between types in a given set, or an estimate of the Bayesian classification of the object type), as opposed to selecting the point for labeling randomly from the unlabeled set. We employed the modAL package to choose objects using a selection criterion based on which objects are most ambiguous, or in other words, objects for which the maximum classification probabilities are the lowest. In doing so, we can leverage human/manual classifications where our random forest classifier might be most uncertain. This criterion for selecting objects for manual classification is known as “uncertainty sampling”. Figure 9 shows a diagram of the process.
From the initial random forest classification, each glitch has four probabilities related to the four possible object types: PS, PS+, CR, and EG. The most likely (highest) probabilities for all glitches are ranked, and the uncertainty sampler takes the lowest value of those probabilities. In other words, the uncertainty sampler selects a glitch that the random forest classifier is the least confident in (or most uncertain about) to manually label and therefore add to the active learning training set 30. This is shown in the bottom right-hand section of Figure 9.
To demonstrate the effects of active learning, we tested the algorithm on a set of 851 labeled glitches from 2021 at 90 GHz and 150 GHz for detector arrays PA4 and PA5 (of which 63 are PS, 44 PS+, 363 CR, and 381 EG). We begin with a training set composed of 2280 glitches from 2019 at 90 GHz in PA5. We added 300 training points using two methods, active learning and then random selection, that were chosen from a set of 2555 glitches from 2021 at 90 GHz and 150 GHz in PA4 and PA5. It is important to note that the test set and additional training points are comprised of data from different arrays, frequencies, and years which could have different noise properties or detector performance than the original training set with only 2019 at 90 GHz from PA5 data. The test set remained the same throughout, in order to ensure that the accuracy was not increasing due to removing difficult-to-classify test points.
The results from this can be seen in Figure 10 where the left plot shows the confusion matrix that arises from classifications based on the initial training set. The middle and right plots show the confusion matrices where additional training points have been chosen though active learning (middle) or random selection (right). In both cases, the overall accuracy of classification increases modestly, by and percentage points, respectively. We note that the classification accuracy for the PS category increased significantly, by close to 14 percentage points, when using an active learning adjustment and by only 5 percentage points when supplementing the training set randomly.
Similarly, the accuracy for PS+ increased by 11 percentage points when utilizing active learning, as opposed to 5 percentage points with random selection. This is promising: as the volume of data increases with successive observing seasons, new detector arrays, and looking to future telescopes such as SO, we are now motivated to use active learning to decrease the amount of manual labeling needed in a training set. Active learning also allows us to ensure that we are choosing test points that are difficult for our algorithm to classify, and therefore gaining a better understanding of its performance.
6 Stellar Flare Simulations
To test the classification algorithm on a point source with known position and amplitude, we inject simulated signals of stellar flares into real TODs. This allows us to test our classification pipeline on timestreams that contain real noise plus the addition of signals from simulated sources. We then use the known pointing matrix, which describes the sky coordinates of the telescope pointing as a function of time, to project an instrumental beam profile, as is appropriate for a point source, into the TOD timestreams. This process projects the 2D spatial beam into the timestreams for all detectors that scan over the simulated source.
We vary the amplitude of the injected source over time as:
where is the initial amplitude of the flare, is the half-life of the decaying flare, and is the start time of the flare, which we select to be when the source enters the focal plane for the first time in the TOD. Stellar flares can display a variety of behaviors across the electromagnetic spectrum and can be fit with different profiles. For instance, MacGregor et al. 32 fit two millimeter flares detected by the Atacama Large Millimeter/submillimeter Array 50 ALMA with Gaussian profiles. Salter et al. 42, on the other hand, used an exponential curve to fit the decaying millimeter flux of the equal-mass, highly eccentric, close binary system DQ Tau observed with the Institute for Radio Astronomy in the Millimetre Range Plateau de Bure Interferometer 21 IRAM PdBI, the Combined Array for Research in Millimeter-wave Astronomy 11 CARMA, and the Submillimeter Array 27 SMA. In this paper, we opt for a simple model in which an instantaneous jump in flux is followed by an exponential decrease.
Figure 11 shows an example of a stellar flare injected into a TOD. The simulations are run for sources with an initial amplitude of 12,000 and mJy and flare profiles with half-lives of seconds. These amplitudes were chosen as this specific TOD has an average noise level of mJy, and we wanted to test a case where most detectors would be well above the cut-off and a case closer to the limit. These half-lives were chosen to cover a wide range of flare durations. The shortest ( s) half-life reproduces a situation where the flare decays significantly within a single scan, whereas the longest ( s) half-life was chosen to maximize the number of detections within the 11-minute TOD. The TODs are loaded in raw data acquisition units (DAU) then converted into mJy, at which point the source signal is injected. The TOD is then converted to K to be compatible with TOD loading and cuts packages, at which point it can be run through our pipeline.
7 Results and Discussion
7.1 Classification Results
We tested our trained forest on a set of real data to determine how effective it is at classifying true snippets, all labeled as outlined in Section 5.1. Our training set includes 2580 glitches which consist of 272 PS, 167 PS+, 1119 CR and 1,022 EG objects respectively. Of these glitches, 2280 are from 2019, 90 GHz from PA5, and the remaining 300 are from 2021, 90 GHz and 150 GHz from PA4 and PA5. For testing, we had 3865 glitches: 282 PS, 207 PS+, 1672 CR, and 1704 EG. This consisted of 760 glitches from 2019, 90 GHz, and PA5 as well as 3,105 from 2021, 90 GHz and 150 GHz from PA4 and PA5. Figure 12 shows the resulting confusion matrix, where each value indicates the decimal percentage of each glitch category on the -axis being predicted as the category shown on the -axis. For perfect performance, the resulting confusion matrix would be an identity matrix. Our overall accuracy is 90%, and notably, we achieved a 94% accuracy for PS objects. For CR objects, are labeled EG and of the EG objects are being labeled as CR. This is expected as some electronic glitches, such as large-area heat-ups of the focal plane, look similar to cosmic rays and vice versa.
To test the performance of the classifier, we also examined the precision, recall, and F-1 score (a combination of precision and recall) for each of the classification categories. As illustrated in Equation 5, precision is the proportion of glitches predicted to be within a category that were predicted correctly, while recall measures how many glitches of a certain category were correctly predicted. A high precision score therefore indicates a low false positive rate (few objects incorrectly identified), whereas a high recall indicates a low false negative rate (few objects missed). The F-1 score is a harmonic mean of the precision and recall of the forest. They are given by:
where tp is the number of true positives, fp is the number of false positives, fn is the number of false negatives, is the precision, is the recall, and F-1 is the F-1 score. The results of our test set can be seen in Table 1. All the F-1 scores across the four categories are greater than 89%, which indicates a well-performing model. We consider a well-performing model to be one where the metrics discussed above are ; however, we also favor as high a value of precision as possible, as false detections of point sources that are left in the TODs will have a negative impact on the maps.
Glitch Category | Number of Snippets | Precision [%] | Recall [%] | F-1 Score [%] |
PS | 282 | 96.0 | 93.6 | 94.8 |
PS+ | 207 | 96.6 | 82.6 | 89.1 |
CR | 1672 | 87.2 | 92.6 | 89.8 |
EG | 1704 | 91.3 | 87.7 | 89.4 |
We consider a classification of an individual object with a true underlying type to be “high-probability” if it is predicted to be a particular type with a probability, defined as the percentage of decision trees in the forest that predict the glitch to be a given type, to be greater than 70%. After filtering for high-probability classifications in our test set, we are left with six PS+ objects that were mistakenly classified as PS objects, and examine these individually to understand the misclassifications. The TODs and focal plane positions for these snippets are shown in Figure 13. In these TODs, another glitch appears alongside a point source, but the detectors affected by the other glitch are located near to the PS in the focal plane. This resulted in the “Normalized Maximum within 0.1” statistic for all cases to be equal to one. Similarly, the “Normalized Maximum” values for these objects range from 0.27 to 0.51 and the “ and Extent Ratios” range from 0.06 to 0.24, which is consistent with values observed from objects with the PS designation rather than PS+. As described in Section 4, the three statistics mentioned above are usually the only ones that separate cleanly between the PS and the PS+ categories, while the other statistics used for classification are consistent with expected values for both types. In this case, the fact that the statistics we input as the features of our algorithm suggest a PS categorization (as is shown in Figure 2) is the dominant reason for the misclassification. We repeat that these three statistics are important for the classification results, as can be seen in Figure 7. Future work should be done to determine another statistic to aid classification in these instances. One possibility would be a statistic that identifies spikes in the signal that are beam-shaped as we would expect from a source but not from a cosmic ray or electronic glitch. Note that human bias also affects the results, as our training and testing labels are from human labeling.
7.2 Improved Transient Search Maps
In addition to removing glitches from the TODs prior to map-making, our algorithm is also able to classify astrophysically interesting bright sources from the TODs and generate a list of detection times and locations of these sources. This represents a complementary method to map-based transient search algorithms e.g., those presented in 12. For map-based transient searches, “depth-1” maps, or maps from a single array and frequency in which each point in the map drifts only once through the focal plane—roughly speaking, a single observation session—are used. Biermann et al. 12 provides details on the ACT depth-1 maps and the results of blind transient searches on these maps. Current map-based searches have focused on finding faint sources outside the Galactic plane. This is partially due to the limitations of identifying many sources that are spatially close together in the maps, as occurs within the Galactic plane. Because our algorithm processes each TOD snippet individually, we do not, however, encounter this limitation to the same extent as map-based searches.
To demonstrate that our algorithm can prevent genuine sources from being removed from depth-1 maps, we modify the ACT cuts software to not remove any source which has a probability of being a PS greater than 70%, as determined by the random forest, and compare the results to what is obtained with the original cuts software, doing so on a series of stellar flare simulations described in Section 6. The steps of this modified algorithm can be seen in Figure 14, where the existing portions of the pipeline that we modified are shown in purple and our new glitch classification algorithm is shown in teal. In order to demonstrate the efficacy of our method, we generate a bright, relatively stable source that stays at by injecting a stellar flare with a peak amplitude of 12,000 mJy and a half-life of 5000 s. The results can be seen in Figure 15. We note that the original algorithm removes the source from the map, whereas our modified algorithm restores the source. A ringing pattern along the scan direction through the source is visible in this map cut-out. This is a form of subpixel model error bias (related to regression dilution) caused by treating the sky as being stepwise constant in each pixel despite the true signal changing smoothly, and is unrelated to our cuts implementation 38. Additionally, as we are only producing maps for the purposes of source detection, there could be a lack of convergence in the maps which could also produce some of this affect.
To test the ability of our algorithm to identify sources with different properties, we vary the initial peak amplitude and half-life of the simulation. Recall that the cuts algorithm identifies glitches by locating spikes in the TOD with an ; the TOD used in this simulation has an average noise of mJy, though individual detectors can have more or less noise than this value. We thus start our exploration with an amplitude of 12,000 mJy to ensure that we are comfortably above the threshold. The top row and bottom row per amplitude of Figure 16 shows the maps with each of the original and modified cuts pipelines, respectively. For this large amplitude, the source was at least partially cut out of the maps using the original pipeline, but completely restored with the new pipeline for all simulated source half-lives except 0.5 s, which was classified as a PS, but only with a probability of 70% (which is at the threshold of what we consider a high-probability detection). In the case of the 6000 mJy amplitude source, it is above the noise value for most, but not all, of the detectors. In this case, the source is classified as a PS for and 5 s but only with a probability of 68%, which is below our detection threshold. The source appears in the map made with the original algorithm and also in our modified algorithm for and 250 s and the center of the source is fully restored for and 750 s which is necessary for map-based analysis of the source properties. The reason the source can appear in the original maps even though some detectors are cut is because a portion of the detectors see the source below the threshold, while the ACT cuts algorithm requires that the source has an for a detector to be cut. There is some utility, however, to being able to identify a source that is only partially detected/cut, as one can measure the amplitude variations between snippets to aid classification of the source when it is later detected in a map-based search. Partial identification of sources also generates a list of times and locations of observed sources that are useful for the map-based search methods described above. Note that we chose the probability threshold in order to reduce the number of false positives for map-making, but for transients searches one could lower this threshold and the source would have been recovered for both amplitudes with a half-life of 0.5 s.
Finally, we tested our algorithm for an amplitude of 3000 mJy, which is similar to the amplitude of the brightest transient found in ACT ( mJy; Biermann et al. 12, Li et al. 31). We note, however, that previous transient searches have excluded the galactic plane, where we have found many sources with higher amplitudes that we have successfully identified and restored with the new algorithm. The injected source was only identified for and 750 s with all other values of half-life resulting in the source remaining below the detection threshold. We would like to understand if this failure to correctly classify the injected source is due to a limitation of our classification performance or a result of the earlier step in the process where the snippet is flagged by the cuts algorithm.
To do so we ran the simulated sources through a modified cuts algorithm that we call the “ideal cuts” algorithm, that flags every detector that observes the source with an amplitude greater than 500 mJy, rather than the usual approach used for ACT analysis which requires the signal to be above the threshold. Given that this is a simulation and hence we know a source is present, we can isolate the “detection” versus the “classification” parts of our analysis, removing any uncertainty in the detector cuts pipeline and focusing only on the performance of the classification algorithm presented in this paper. Employing this procedure (where a source with an amplitude greater than 500 mJy detected by any detector is flagged), we were able for each simulated amplitude to detect the injected sources for all values of the half-lives and initial amplitudes we simulated. We provide more details on this test in Appendix Section A. The failure to detect the source by the detector cuts algorithm is due to the source having an below the threshold. Due to the importance of summary statistics related to the focal plane statistics in our classification algorithm (as is shown in Figure 7), we are able to classify the sources, provided they are detected by our “ideal cuts” algorithm even though the source signal is very faint in the TODs. The cutoff is a known limitation of the cuts algorithm, highlighting the need for future work to create an improved cuts algorithm that would enable our classification algorithm to find lower-amplitude sources. An improved cuts algorithm would enable TOD based classification algorithms, such as the one outlined in this paper, to identify dim sources prior to map-making. This would not only aid in quick identification of possibly interesting sources, but also provide a list of candidate positions and times to search for sources in the depth-1 maps.
7.3 Point Source Variability
If a source is observed multiple times, as the focal plane scans back and forth in azimuth at constant elevation, we can measure changes in the amplitude of the source on scales of less than a minute for an ACT-like experiment. This has important implications for transient science, where some sources, like flaring stars, can vary on time scales of minutes or seconds, whereas other sources, such as extragalactic transients like GRBs, TDEs and SNs, vary on much longer time scales of days or more. To issue timely and accurate alerts of the detection of such events to the scientific community, we need our code to be able to measure changes in amplitude on similar timescales. Here, we demonstrate the capabilities of our algorithm to retrieve variability information on classified point sources using the stellar flare simulations described in Section 6.
We model the amplitude of a source in our timestreams with a non-linear least-squares fit to the following model
where is the amplitude measured by a detector observing the sky at right ascension/declination position , is the intrinsic amplitude of the source located at position , is the azimuthally-averaged, normalized beam profile of the telescope, and is an offset to account for the baseline of the signal being nonzero. The independent variable in the fit is and the fit parameters are , and . Before performing the fit, we create a copy of the snippet with the peaks excised, and we subtract this copy from the original snippet, leaving only the peaks generated by the point source in the data. This removes any remaining large-scale drift across the length of the snippet, which we find significantly improves the fit. We then discard half of the 0.5 s buffer from each of the beginning and end of the snippet, i.e., the first and last 100 samples to reduce the amount of noise that goes into the fit, as we know that the snippets have a buffer of 200 samples on either side of the source signal. Finally, we obtain initial guesses for the non-linear fit by averaging the peak amplitude values of all detectors, for the initial value of , and the detector positions at their peak values, for ; is initially set to zero. We then run the curve_fit non-linear fitter as implemented by SciPy.
We group snippets that are classified as high-probability point sources within 0.01 of each other (corresponding to the standard error on our point source localization in the timestreams), treating them as the same astronomical object, and then we do the fit described above for each scan of the array across the source. This produces a light curve with one point per scan. Our best-fit light curves to simulated stellar flares injected into real data are shown in Figure 17. The position of the source relative to the scan pattern differs for each TOD, causing the time intervals between the source showing up in successive scans to vary from TOD to TOD. This can be seen in the different segment lengths of the individual best-fit light curves (dashed lines) in Figure 17. For this reason, we linearly interpolated the best-fit light curves into 10 s time bins. The amplitudes were then averaged in each time bin to obtain a mean and standard deviation, shown with the shaded region in Figure 17. We only simulate stellar flares with half-lives above 25 seconds because shorter flares decay too quickly to be detected for more than one scan.
To evaluate the level of confidence with which we can detect short-term variations with this method, we calculate the likelihood ratio (or , assuming a Gaussian likelihood) of our best-fit light curve from a single TOD to a flat line at the average flux value; the latter is the null hypothesis of no variation in the source. The likelihood ratio can be converted to a p-value for the null hypothesis, or equivalently a significance of the measurement of variability.333The conversion of the likelihood ratio to a -value assumes that the former has a distribution, which is true for numerous samples; in our case the number of samples is low, so the values we report should be taken as approximate. For uncertainties in our measurement, we use the standard deviation from fits to 11 different TODs into which the same simulated flare is injected. The results of this analysis are presented in Table 2.
Variability Significance [] | ||
Half-life [s] | 12,000 mJy | 6000 mJy |
750 | 2.8 | <1 |
500 | 4.2 | 1.7 |
250 | 5.8 | 2.0 |
100 | 10.9 | 1.6 |
25 | 5.3 | <1 |
Although the algorithm’s ability to detect a varying signal is currently limited to very bright objects, due to the threshold of the cuts algorithm (as discussed in the previous section), these results highlight the potential of the algorithm to detect quickly decaying flares, sub-minute flares with at least two observations, in the timestreams. This will be an essential tool in detecting fast varying (sub-minute to minute) transients.
8 Conclusions
Our improved data cuts algorithm enables classifications of glitches into point sources, point sources with another coincident glitch, cosmic rays, and electronic glitches from the output of the ACT glitch finder. We have an overall accuracy of 90%, with 94% of PS objects being correctly classified. We are able to restore high-amplitude sources that previously would have been removed from the maps and flag the sources that would have partially passed the cuts. This enables us not only to ensure that sources are not removed from the maps but also to generate a list of times and positions of transients. Such a list can be used to issue rapid alerts and also provide informative priors for map-based searches.
The algorithm selects sources that appear multiple times in a TOD. The amplitudes of such sources can be computed across scans to generate sub-minute to minute timescale light curves. We tested the recovery of simulated light curves of the stellar flares using this pipeline. TOD-specific systematics were averaged out in order to better understand the performance we could generally expect from the amplitude computation. We found that for a bright 12,000 mJy source we can detect variability with for a flare with half-life s, and for a dimmer, 6000 mJy source the significance was 1.5–2 for half-lives between 100 s and 500 s, with no detection of longer or shorter flares. These results are from ACT PA4 and PA5, and with more sensitive instruments, like the upcoming SO 3, 2, the ability to determine whether a source is varying on short time scales will improve. Finally, even when a candidate transient has low SNR, or is only seen in one or two scans, we can perform forced-aperture photometry on the source location directly in the TOD to obtain a longer light curve with a baseline flux before the transient appears.
We are adapting our algorithm for SO. In addition to providing data cuts for the main map-making pipeline, its transient detection capabilities are highly complementary to the map-based transient searches which are being developed. In addition to the benefits mentioned above—preventing transients from being excised from the data and providing a list of transient candidates to map-based searches—it can create high time-resolution light curves (sub-minute to minute) of transients, which cannot be readily extracted from maps. This will be helpful in issuing faster transient alerts to the broader community, since current ACT map-based transient searches require multiple TODs and the creation of depth-1 maps. Additionally, the light curves retrieved using this algorithm could allow us to rapidly classify the potential transient type (e.g. stellar flares, which are short duration, versus extragalactic explosions like gamma ray bursts, which are long duration). Furthermore, our experience thus far indicates that TOD searches may perform better on bright transients in the Galactic plane where the source density is higher, and it is more challenging for map-based searches to disentangle nearby sources from each other.
Other avenues for future work include, for instance, adding more statistics to reduce misclassified glitches, particularly for a point source with another spatially coincident glitch in the focal plane. Another interesting avenue would be to create a finer electronic glitch categorization to aid in the understanding of telescope systematics. This would especially help diagnose issues that arise when building a new telescope or implementing new hardware. This could be achieved by first implementing an unsupervised clustering algorithm to determine which categories of electronic glitches arise. We focused on using summary statistics in order to have a light-weight algorithm that can be run in real time during pre-processing. Other approaches, even though they are more memory intensive, such as image classification on the focal plane and timestreams, should also be explored. Another important area for future development is to improve the initial glitch detection algorithm so that it has a lower cut-off, since the relatively high threshold currently being used is a limiting factor in our performance. We are currently developing methods using machine learning to better detect glitches and sources in the data, with the goal of integrating detection into the classification algorithm. Finally, we have demonstrated our ability to recover sub-minute light curves directly from the TODs, but our simple fitting method could be further improved, particularly to obtain better error estimates for the light curves.
Acknowledgements
Support for ACT was through the U.S. National Science Foundation through awards AST-0408698, AST-0965625, and AST-1440226 for the ACT project, as well as awards PHY-0355328, PHY-0855887 and PHY-1214379. Funding was also provided by Princeton University, the University of Pennsylvania, and a Canada Foundation for Innovation (CFI) award to UBC. ACT operated in the Parque Astronómico Atacama in northern Chile under the auspices of the Agencia Nacional de Investigación y Desarrollo (ANID). The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. Detector research at NIST was supported by the NIST Innovations in Measurement Science program. Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund–Research Excellence, and the University of Toronto. We thank the Republic of Chile for hosting ACT in the northern Atacama, and the local indigenous Licanantay communities whom we follow in observing and learning from the night sky. YG acknowledges support from the University of Toronto’s Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a program of Schmidt Sciences. ADH acknowledges support from the Sutton Family Chair in Science, Christianity and Cultures, from the Faculty of Arts and Science, University of Toronto, and from the Natural Sciences and Engineering Research Council of Canada (NSERC) [RGPIN-2023-05014, DGECR-2023-00180]. CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through Basal project FB210003. RH acknowledges support from the NSERC Discovery Grant Program RGPIN-2018-05750 and the Connaught Fund. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. The authors at the University of Toronto acknowledge that the land on which the University of Toronto is built is the traditional territory of the Haudenosaunee and, most recently, the territory of the Mississaugas of the New Credit First Nation. They are grateful to have the opportunity to work in the community on this territory.
Appendix A Classification Results with an Ideal Cuts Algorithm
The current cuts algorithm requires an cutoff of 10 which results in only bright sources to be flagged. In order to test if our algorithm would be able to classify dimmer sources if they were flagged by the cuts, we utilized the stellar flare simulations to create “ideal cuts”. Since the sources are simulated, we know the amplitudes at all times. We create modified cuts objects that flag every detector at a given time that has an amplitude greater than 500 mJy, chosen to be just below the average noise of the detectors. We ran these modified cuts through our classification algorithm, and the results can be seen in Figure 18. For all half-lives and amplitudes, we detected the source using these ideal cuts. This is very promising and is likely due to the importance of the focal plane statistics to our classification algorithm, as we see Figure 7. In general, we see a decrease in probability for later scans as the source’s amplitude is decaying following Equation 6. However, this trend is not strictly followed for the first couple scans, as the ACT focal plane is a hexagon oriented such that the scan begins at one of its vertices. Thus, when the source first enters the focal plane, it only hits a few detectors. As the sky rotates through the focal plane, which is scanning at constant elevation, the source moves towards its center and hits more detectors; finally, as it exits the focal plane it again only hits a few detectors. We note that for 3000 mJy, fewer than four detectors were flagged for the first three scans and thus were not classified, as we require 4 detectors for classification.